OpenStudy (anonymous):

A(2,-2) B(11,-4) find AC + BC is a bare minimum on X axis... Show ur work too mates

OpenStudy (anonymous):

can't wait to see the answer to this...

OpenStudy (solomonzelman):

more information? what course is this from?

OpenStudy (solomonzelman):

as it is now, I don't think there is an answer, and that's what sate is most likely saying to us-:(

OpenStudy (solomonzelman):

or wait...

OpenStudy (solomonzelman):

okay, say you have a function \(\large\color{black}{ y={\rm AC} +{\rm BC}}\) \(\large\color{black}{ y={\rm C(A} +{\rm B)}}\) I will assume to exclude any values of C equal to or less than 0. I get that for \(\large\color{black}{ y\rightarrow0}\), \(\large\color{black}{ C\rightarrow0}\). this is the best I can get. (assuming that A and B are positive)

OpenStudy (solomonzelman):

with more info, we could do more

OpenStudy (anonymous):

that is all the info ive obtained

OpenStudy (anonymous):

oh and the course is Geometry

OpenStudy (anonymous):

honors Geometry

OpenStudy (anonymous):

@SolomonZelman

jimthompson5910 (jim_thompson5910):

check out this page http://hom.wikidot.com/heron

OpenStudy (anonymous):

Thx Jem @jim_thompson5910

OpenStudy (kamibug):

\(\huge\cal\color{indigo}{Welcome~to~Openstudy!~:)}\) @Da_Homie

OpenStudy (solomonzelman):

oh, tnx for editing it.

OpenStudy (solomonzelman):

\(\large\color{black}{ A(2,-2)~~~B(11,-4) }\) and you may be mean, ` find C so that AC + BC is a bare minimum on X axis` ?

OpenStudy (anonymous):

yea

OpenStudy (solomonzelman):

oh, that is much better. I am not very sure what a bear minimum on x-axis means.

OpenStudy (anonymous):

lol what is the minimum distance it can be

OpenStudy (anonymous):

minimum distance Ac + BC is

OpenStudy (solomonzelman):

for it to be on the x-axis. right?

OpenStudy (solomonzelman):

for it, for C to be on x-axis. \(\large\color{black}{ A(2,-2)~~~B(11,-4) }\) https://www.desmos.com/calculator/wyfre5igq8 showing the pic of the 2 points.

OpenStudy (anonymous):

yes exactly my friend

OpenStudy (solomonzelman):

it would be logical to say that the closer 2 sides to each other are, the more likely will it be a minimum: (As, \(\large\color{black}{ 4^2+4^2<8^2+2^2 }\) . not exactly like this, but idea is same) \(\large\color{black}{ AB=\sqrt{9^2+4^2}=\sqrt{81+16}=\sqrt{95} }\) (just in case we need it)

OpenStudy (solomonzelman):

it would be logical to say that the closer 2 sides to each other are, the more likely will it be a minimum: (As, \(\large\color{black}{ 4^2+4^2<8^2+2^2 }\) . not exactly like this, but idea is same) \(\large\color{black}{ AB=\sqrt{9^2+4^2}=\sqrt{81+16}=\sqrt{95} }\) (just in case we need it)

OpenStudy (solomonzelman):

when \(\large\color{black}{ AC=BC }\) would get a minimum distance: Solving for C. \(\large\color{black}{ \sqrt{(2-x_c)^2+(-2-y_c)^2}=\sqrt{(11-x_c)^2+(-4-y_c)^2} }\)

OpenStudy (solomonzelman):

where point C is \(\large\color{black}{ (x_c,y_c) }\)

OpenStudy (anonymous):

thats the equation???

OpenStudy (solomonzelman):

I will use just x and y, without sub c.

OpenStudy (solomonzelman):

got disconnected, dang it

OpenStudy (solomonzelman):

I am raising both sides to second power, and plugging in y=0, since that is what I want to find (at the y axis)

OpenStudy (anonymous):

ok, im with u

OpenStudy (solomonzelman):

\(\large\color{black}{ (2-x)^2+4=(11-x)^2+16 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ 4-4x+x^2+4=121-22x+x^2+16 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ 4x+x^2+4=117-22x+x^2+16 }\) \(\large\color{black}{ 4-4x+4=121-22x+16 }\) \(\large\color{black}{ -4x+8=137-22x }\) \(\large\color{black}{ 16x=129 }\) \(\large\color{black}{ x=129/16 }\)

OpenStudy (anonymous):

i got 8.0825 :( i think its wrong

OpenStudy (solomonzelman):

So I am thinking it would be \(\large\color{black}{ C~~(129,16~,0) }\)

OpenStudy (solomonzelman):

when I graph the points I get: https://www.desmos.com/calculator/ujh6gbdujs

OpenStudy (anonymous):

what did you get for X??

OpenStudy (solomonzelman):

the x coordinate of the point I got is 129/16

OpenStudy (anonymous):

but that would be a decimal right?

OpenStudy (anonymous):

8.0825

OpenStudy (solomonzelman):

8.0625 is what I get

OpenStudy (anonymous):

8.0625*

OpenStudy (anonymous):

i meant 6 lol

OpenStudy (solomonzelman):

I am not 100% sure about my solution though.

OpenStudy (anonymous):

same

OpenStudy (solomonzelman):

looks like I got it write, like after graphing, doesn't seem to have some crazy non sense

OpenStudy (anonymous):

how did u graph that anyway??

OpenStudy (solomonzelman):

I polted in a graphing calculator

OpenStudy (solomonzelman):

@plotted

OpenStudy (anonymous):

ohh nvm i was thinkng something else lol

jimthompson5910 (jim_thompson5910):

Step 1) Plot the two points A and B Step 2) Pick a point you want to reflect over the x axis, say B. When you reflect B over, it will land on B' Step 3) Construct the line AB' Step 4) The point of intersection between AB' and the x axis is where C should go to minimize AC+BC. This is shown in the link http://hom.wikidot.com/heron